# TACO: Topics in Algorithmic COde generation dataset

### v2

@@ Abstract
We introduce TACO, an open-source, large-scale code generation dataset, with a focus on the optics of algorithms, designed to provide a more challenging training dataset and evaluation benchmark in the field of code generation models. TACO includes competition-level programming questions that are more challenging, to enhance or evaluate problem understanding and reasoning abilities in real-world programming scenarios. There are 25433 and 1000 coding problems in training and test set, as well as up to 1.55 million diverse solution answers. Moreover, each TACO problem includes several fine-grained labels such as task topics, algorithms, programming skills, and difficulty levels, providing a more precise reference for the training and evaluation of code generation models. The dataset and evaluation scripts are available on Hugging Face Hub111https://huggingface.co/datasets/BAAI/TACO and Github222https://github.com/FlagOpen/TACO.
@@ 3 Dataset Construction
The development of the TACO dataset is meticulously divided into two distinct phases: data collection and data processing. In the data collection phase, we focused on gathering raw data primarily from a diverse range of programming competition platforms. This phase was further enriched by integrating additional open-source datasets, thereby expanding the dataset’s scope and variety. Following the collection, the data processing phase commenced, involving a series of intricate steps. Initially, this phase entailed code de-commenting and de-duplication to ensure data quality and uniqueness. Subsequently, we augmented the test samples to enhance the dataset’s robustness. In this phase, we also engaged in the classification and generalization of tasks and skill sets, using insights gained from a variety of programming contest websites and open-source code generation datasets. This comprehensive approach enabled us to establish a well-defined and detailed format schema for the TACO dataset.
@@ 3.1 Data collection phase
During the assembly of the TACO dataset, we conducted an exhaustive review of various programming contest platforms as well as existing code generation datasets. Owing to legal restrictions and the intricate nature of platform-specific API interfaces, our primary data sources were platforms such as CodeChef, CodeForces, HackerRank, and GeeksforGeeks. We also integrated pre-existing datasets, including APPS, CodeContest, and Description2code. To construct and optimize the TACO dataset, we employed two primary methodologies: First, we utilized web scraping techniques to extract problem information from programming competition websites; second, we augmented existing open-source datasets with algorithmic skill labels and other relevant tags.
@@
To guarantee the dataset’s integrity and the comprehensiveness of its tagging metadata, we devised bespoke HTML parsers tailored to each of the four major programming contest platforms: CodeChef, CodeForces, HackerRank, and GeeksforGeeks. These parsers are engineered to accommodate the diverse syntactic layouts present in the problem texts of their source websites. Throughout the parser development phase, we confirmed the accuracy of each specialized parser via extensive manual sampling. On exceptional occasions, we also employed OCR tool APIs (e.g., SimplePix) to facilitate the conversion of svg-formatted information into LaTeX format. For a detailed discussion of validation procedures and data scraping techniques used for these platforms, refer to appendix A. Additionally, with regard to the algorithmic labeling designed to augment the APPS and CodeContest datasets, our HTML parser similarly crawls algorithmic skill labels from sources such as LeetCode, Kattis, CodeWars, and CodeForces.
@@
Upon completion of the TACO dataset assembly, we undertook a rigorous analysis to assess the degree of overlap between TACO, APPS, and CodeContest in terms of problem quantity. Initially, we executed problem correspondences predicated upon the URLs corresponding to each problem. Subsequently, we expanded our correspondence framework to include plaintext analyses, which entailed the removal of whitespace and newline characters. Our computations revealed that there are 3,392 duplicate problems shared between APPS and CodeContest. The degree of problem overlap between TACO and CodeContest stands at 0.4617, while the overlap coefficient with APPS is 0.3778. As depicted in Figure 2, 29.5% of the problems constituting the TACO dataset are sourced from our recent web crawling endeavors; among the remaining problems, 24.3% are exclusive to the APPS dataset, 32.7% are unique to CodeContest, and the residual 13.4% are featured in both APPS and CodeContest.
@@
As detailed in Table 1, the TACO dataset encompasses data from multiple programming contest platforms, listing the diverse data sources corresponding to each website. For websites with multiple sources of data, we have executed data merging and de-duplication procedures. It is noteworthy that certain platforms, such as Aizu, AtCoder, and HackerEarth, already feature data integrated from CodeContests and Description2code, and include the requisite algorithmic tags; therefore, additional crawling efforts were deemed unnecessary for these sites.
@@ 3.2.2 Functional processing
Code De-duplication: Given that our data collection draws from solutions submitted by multiple users and integrates various open-source datasets, the propensity for encountering similar or duplicate code within individual problems is elevated. To mitigate this issue, we employed a targeted de-duplication strategy. Specifically, we utilized the MinHash algorithm in conjunction with the Jaccard similarity index to achieve near-deduplication. A Jaccard similarity threshold was established at 0.85, and MinHash-assisted clusters of duplicate code were generated. Subsequently, these clusters were condensed into unique code files based on the Jaccard similarity metric.
Code De-commenting: In the code submissions retrieved from programming contest platforms, such as CodeChef, we observed a significant prevalence of comments. These comments, while ostensibly redundant, not only inflate the code’s length but also introduce potential noise into subsequent analyses. To address this issue, we employed the Abstract Syntax Tree (AST) parsing methodology to excise these comments. Specifically, the collected code is initially converted to AST format, whereupon comment nodes are identified and eradicated. Following this, the AST structure is reverted to its original code format. It is noteworthy that the AST parsing technique also affords us the capability to execute additional code formatting tasks, such as the elimination of superfluous whitespace.
Supplementation of Test Cases Owing to the unavailability of hidden unit tests on programming websites for crawling, the majority of problems within the APPS dataset are restricted to one or two sets of unit tests as provided in the problem description. This constraint not only undermines the precise evaluation of code correctness but also risks the generation of incorrect code that may pass these limited unit tests. According to AlphaCode, the APPS dataset exhibited a False Positive (FP) Rate of up to 60% when assessed with an average of 20 unit tests, thereby signaling a lack of stringent evaluation criteria. To rectify this limitation, we augmented the code samples with additional unit tests, elevating the average number of unit tests to over 200. AlphaCode’s research indicates that the FP Rate can be substantially reduced to approximately 4% when the average number of unit tests surpasses 200. Specifically, we leveraged OpenAI’s GPT-4 API to generate the input components of unit test pairs. Subsequently, 30 verified correct code samples were executed on these inputs to generate corresponding outputs, which were then scrutinized for consistency. Input-output pairs yielding consistent outputs were deemed valid unit tests. This iterative process was carried out multiple times to ensure that the total number of unit tests reached a minimum threshold of 200.
@@ 3.2.3 Skill algorithm categorization
In the Source Programming Problems dataset, we retained 968 category original labels spanning various domains. These labels were manually annotated by domain experts and offer the most granular level of informational categorization. Specifically, these labels may encapsulate either the subject matter of the problem (e.g., mathematics, geometry, graphs, or strings) or feasible solution strategies (e.g., dynamic programming, brute-force approaches, or binary search). It is noteworthy that a single problem may bear multiple such labels concurrently. For a comprehensive view of the original label distribution across all problems in the dataset, refer to Figure 3. As evidenced by this figure, the distribution of these foundational labels is highly irregular and diverse. Consequently, it is crucial to systematically categorize and generalize these algorithmic labels.
A usability analysis was undertaken to elucidate both the quantity and distribution of raw labels. Our findings suggest that an excessive and intricate array of algorithmic labels could potentially impede the model’s training and inference processes. Consequently, we executed a thorough generalization of the initial set of 968 category algorithmic labels. Guided by three key considerations, we ultimately reclassified all labels into 36 discrete categories:
@@
Label Quality: Following meticulous organization and generalization, the resulting label set should accurately and profoundly capture the quintessential attributes of programming problems and their corresponding solutions.
@@ 5.1 Experimental Setup
Similar to mainstream code generation evaluations, for each programming problem, we allow the model to generate 200 attempts, using pass@k (k=1, 10, 100) as the evaluation metric. Therefore, during the model evaluation process, it is necessary to set the hyperparameters of the generation model. Through extensive generation experiments, we found that the model is highly sensitive to the difficulty of the problem. For more challenging problems, the model requires a more diverse range of generated results to potentially produce code that passes the tests. Consequently in the testing process, we use a setting of 0.95. The temperature settings are differentiated based on problem difficulty levels: we assign a value of 0.2 for easy problems, 0.6 for medium and medium hard problems, and 0.8 for hard and very hard problems.
@@ 5.2 Results
We selected several classic coding models for evaluation, including codellama-7b, codegen25-mono, starcoder-1b, and starcoder-15.5b, and also conducted tests on GPT-4 for comparison. The evaluations were performed on 200 questions across five different difficulty levels, and the results are presented in the Table 5:
@@ 6 Other Application Scenarios
This subsection provides an exhaustive investigation into the versatile utility of the TACO dataset across diverse application domains. The TACO dataset incorporates labels associated with algorithmic skills and is designed to assimilate these skills coherently into code-generation frameworks. Specifically, we examine the dataset’s prospective contributions to several sectors, including code comprehension, educational pedagogy, algorithmic code recommendation, and applications rooted in Language Learning Models (LLMs).
@@
Education and learning: In the educational domain, the TACO dataset holds promise for a diverse range of applications, encompassing the creation of programming textbooks, curriculum development, and the assembly of problem repositories for practice. Learners can enhance their programming proficiencies by tackling the intricate programming challenges contained in the dataset, each annotated with algorithmic skill labels. These labels facilitate a more profound comprehension of the relationship between specific problems and their affiliated algorithms. Moreover, capitalizing on the metadata provided through these labels, educators can streamline the assessment process for student assignments and projects in programming, ensuring adherence to recommended algorithmic methodologies and established best practices.
Code Algorithm Recommendations: The TACO dataset offers promising prospects for contributing to the establishment of an algorithmic recommendation system for coding, which is devised to expedite the process of algorithm selection for software developers. Utilizing the dataset’s array of programming challenges and their associated skill labels, developers can efficiently identify algorithmic exemplars that align with the objectives of their ongoing projects, thereby resolving specific coding conundrums. Furthermore, code editing platforms can exploit the metadata embedded in the TACO dataset to furnish developers with more precise suggestions for code auto-completion, encompassing a broad spectrum of algorithmic and data structural options.
Potential for application on LLM: The TACO dataset holds considerable promise for applications grounded in Large Language Models (LLMs). This dataset serves multiple functions: it can expedite the software development process through the automated generation of task-specific code samples, and it augments the capabilities of LLMs in synthesizing code annotations and documentation to bolster both code comprehensibility and maintainability. Additionally, LLMs that leverage the TACO dataset are capable of addressing queries pertinent to programming and algorithmic topics, thereby furnishing developers and learners with comprehensive and accurate responses. Collectively, these applications contribute to enhancements in programming efficiency, code quality, and the acquisition of programming knowledge, thereby stimulating ongoing advancements in the fields of programming and algorithms.

### v3

@@ Abstract
We introduce TACO, an open-source, large-scale code generation dataset, with a focus on the topics of algorithms, designed to provide a more challenging training dataset and evaluation benchmark in the field of code generation models. TACO includes competition-level programming questions that are more challenging, to enhance or evaluate problem understanding and reasoning abilities in real-world programming scenarios. There are 25433 and 1000 coding problems in training and test set, as well as up to 1.55 million diverse solution answers. Moreover, each TACO problem includes several fine-grained labels such as task topics, algorithms, programming skills, and difficulty levels, providing a more precise reference for the training and evaluation of code generation models. The dataset and evaluation scripts are available on Hugging Face Hub111https://huggingface.co/datasets/BAAI/TACO and Github222https://github.com/FlagOpen/TACO.
@@ 3 Dataset Construction
The development of the TACO dataset is meticulously divided into two distinct phases: data collection and data processing. In the data collection phase, we focused on collecting raw data primarily from a diverse range of programming competition platforms. This phase was further enriched by integrating additional open-source datasets, thereby expanding the dataset’s scope and variety. Following the collection, the data processing phase began, involving a series of intricate steps. Initially, this phase entailed code decommenting and deduplication to ensure data quality and uniqueness. Subsequently, we augmented the test samples to enhance the robustness of the dataset. In this phase, we also engaged in the classification and generalization of tasks and skill sets, using insights gained from a variety of programming contest websites and open source code generation datasets. This comprehensive approach allowed us to establish a well-defined and detailed format schema for the TACO data set.
@@ 3.1 Data collection phase
During the assembly of the TACO dataset, we conducted an exhaustive review of various programming contest platforms as well as existing code generation datasets. Due to legal restrictions and the intricate nature of platform-specific API interfaces, our primary data sources were platforms such as CodeChef, CodeForces, HackerRank, and GeeksforGeeks. We also integrated existing datasets, including APPS, CodeContest, and Description2code. To construct and optimize the TACO dataset, we employed two primary methodologies: First, we utilized web scraping techniques to extract problem information from programming competition websites; second, we augmented existing open-source datasets with algorithmic skill labels and other relevant tags.
@@
To guarantee the dataset’s integrity and the comprehensiveness of its tagging metadata, we devised bespoke HTML parsers tailored to each of the four major programming contest platforms: CodeChef, CodeForces, HackerRank, and GeeksforGeeks. These parsers are engineered to accommodate the diverse syntactic layouts present in the problem texts of their source websites. Throughout the parser development phase, we confirmed the accuracy of each specialized parser via extensive manual sampling. On exceptional occasions, we also employed OCR tool APIs (e.g., SimplePix) to facilitate the conversion of svg-formatted information into LaTeX format. For a detailed discussion of the validation procedures and data scraping techniques used for these platforms, refer to the Appendix A. Additionally, with regard to the algorithmic labeling designed to augment the APPS and CodeContest datasets, our HTML parser similarly crawls algorithmic skill labels from sources such as LeetCode, Kattis, CodeWars, and CodeForces.
@@
Upon completion of the assembly of the TACO dataset, we performed a rigorous analysis to assess the degree of overlap between TACO, APPS, and CodeContest in terms of the quantity of the problem. Initially, we executed problem correspondences predicated on the URLs corresponding to each problem. Subsequently, we expanded our correspondence framework to include plaintext analyses, which entailed the removal of whitespace and newline characters. Our calculations revealed that there are 3,392 duplicate problems shared between APPS and CodeContest. The degree of overlap in the problem between TACO and CodeContest is 0.4617, while the overlap coefficient with APPS is 0.3778. As depicted in Figure 2, 29.5% of the problems constituting the TACO dataset are sourced from our recent web crawling endeavors; among the remaining problems, 24.3% are exclusive to the APPS dataset, 32.7% are unique to CodeContest, and the residual 13.4% are featured in both APPS and CodeContest.
@@
As detailed in Table 1, the TACO dataset includes data from multiple platform programming contests, listing the diverse data sources corresponding to each website. For websites with multiple sources of data, we have executed data merging and deduplication procedures. It is noteworthy that certain platforms, such as Aizu, AtCoder, and HackerEarth, already feature data integrated from CodeContests and Description2code, and include the requisite algorithmic tags; therefore, additional crawling efforts were deemed unnecessary for these sites.
@@ 3.2.2 Functional processing
Code De-duplication: Given that our data collection draws from solutions submitted by multiple users and integrates various open-source datasets, the propensity for encountering similar or duplicate code within individual problems is elevated. To mitigate this issue, we employed a targeted deduplication strategy. Specifically, we utilized the MinHash algorithm in conjunction with the Jaccard similarity index to achieve near-deduplication. A Jaccard similarity threshold was established at 0.85, and MinHash-assisted clusters of duplicate code were generated. Subsequently, these clusters were condensed into unique code files based on the Jaccard similarity metric.
Code De-commenting: In the code submissions retrieved from programming contest platforms, such as CodeChef, we observed a significant prevalence of comments. These comments, while ostensibly redundant, not only inflate the code’s length but also introduce potential noise into subsequent analyses. To address this issue, we employed the Abstract Syntax Tree (AST) parsing methodology to remove these comments. Specifically, the collected code is initially converted to AST format, whereupon comment nodes are identified and eradicated. Following this, the AST structure is reverted to its original code format. It is noteworthy that the AST parsing technique also affords us the capability to execute additional code formatting tasks, such as the elimination of superfluous whitespace.
Supplementation of Test Cases Owing to the unavailability of hidden unit tests on programming websites for crawling, most of the problems within the APPS dataset are restricted to one or two sets of unit tests, as provided in the problem description. This constraint not only undermines the precise evaluation of code correctness but also risks the generation of incorrect code that may pass these limited unit tests. According to AlphaCode, the APPS dataset exhibited a False Positive (FP) Rate of up to 60% when assessed with an average of 20 unit tests, thereby signaling a lack of stringent evaluation criteria. To rectify this limitation, we augmented the code samples with additional unit tests, elevating the average number of unit tests to over 200. AlphaCode’s research indicates that the FP rate can be substantially reduced to approximately 4% when the average number of unit tests exceeds 200. Specifically, we used OpenAI’s GPT-4 API to generate the input components of unit test pairs. Subsequently, 30 verified correct code samples were executed on these inputs to generate corresponding outputs, which were then scrutinized for consistency. Input-output pairs that produced consistent outputs were considered valid unit tests. This iterative process was carried out multiple times to ensure that the total number of unit tests reached a minimum threshold of 200.
@@ 3.2.3 Skill algorithm categorization
In the Source Programming Problems dataset, we retained 968 category original labels spanning various domains. These labels were manually annotated by domain experts and offer the most granular level of informational categorization. Specifically, these labels may encapsulate either the subject matter of the problem (e.g., mathematics, geometry, graphs, or strings) or feasible solution strategies (e.g., dynamic programming, brute-force approaches, or binary search). It is noteworthy that a single problem may have multiple such labels simultaneously. For a comprehensive view of the original label distribution across all problems in the dataset, refer to Figure 3. As evidenced by this figure, the distribution of these foundational labels is highly irregular and diverse. Consequently, it is crucial to systematically categorize and generalize these algorithmic labels.
A usability analysis was undertaken to elucidate both the quantity and distribution of raw labels. Our findings suggest that an excessive and intricate array of algorithmic labels could potentially impede the model’s training and inference processes. Consequently, we conducted a thorough generalization of the initial set of 968 category algorithmic labels. Guided by three key considerations, we ultimately reclassified all labels into 36 discrete categories.
@@
Label Quality: Following meticulous organization and generalization, the resulting set of labels should accurately and profoundly capture the quintessential attributes of programming problems and their corresponding solutions.
@@ 5.1 Experimental Setup
Similar to the evaluations of the generation of standard code, for each programming problem, we allow the model to generate 200 attempts, using pass@k (k=1, 10, 100) as the evaluation metric. Therefore, during the model evaluation process, it is necessary to set the hyperparameters of the generation model. Through extensive generation experiments, we found that the model is highly sensitive to the difficulty of the problem. For more challenging problems, the model requires a more diverse range of generated results to potentially produce code that passes the tests. Consequently in the testing process, we use a setting of 0.95. The temperature settings are differentiated based on problem difficulty levels: we assign a value of 0.2 for easy problems, 0.6 for medium and medium hard problems, and 0.8 for hard and very hard problems.
@@ 5.2 Results
We selected several classic coding models for evaluation, including codellama-7b, codegen25-mono, starcoder-1b and starcoder-15.5b, and we also conducted tests on GPT-4 for comparison. Evaluations were carried out on 200 questions at five different difficulty levels, and the results are presented in Table 5:
@@
6 Limitations
During the construction of the TACO dataset, some programming problems were collected from the APPS and CodeContest datasets and were subsequently reclassified. Therefore, after training code models using the TACO dataset, it is not suitable to use APPS and CodeContest as benchmarks to evaluate the code generation capabilities of the models. For assessing programming challenge capabilities, it is recommended to directly use the test set of the TACO dataset.
@@ 7 Other Application Scenarios
This subsection provides an exhaustive investigation into the versatile utility of the TACO dataset across diverse application domains. The TACO dataset incorporates labels associated with algorithmic skills and is designed to assimilate these skills coherently into code generation frameworks. Specifically, we examine the dataset’s prospective contributions to several sectors, including code comprehension, educational pedagogy, algorithmic code recommendation, and applications rooted in Language Learning Models (LLMs).
@@
Education and learning: In the educational domain, the TACO dataset holds promise for a diverse range of applications, including the creation of programming textbooks, curriculum development, and the assembly of problem repositories for practice. Learners can enhance their programming proficiency by tackling the intricate programming challenges contained in the dataset, each annotated with algorithmic skill labels. These labels facilitate a more profound understanding of the relationship between specific problems and their affiliated algorithms. Moreover, capitalizing on the metadata provided through these labels, educators can streamline the assessment process for student assignments and projects in programming, ensuring adherence to recommended algorithmic methodologies and established best practices.
Code Algorithm Recommendations: The TACO dataset offers promising prospects for contributing to the establishment of an algorithmic recommendation system for coding, which is devised to expedite the process of algorithm selection for software developers. Using the dataset’s array of programming challenges and their associated skill labels, developers can efficiently identify algorithmic exemplars that align with the objectives of their ongoing projects, thereby resolving specific coding conundrums. Furthermore, code editing platforms can exploit the metadata embedded in the TACO dataset to furnish developers with more precise suggestions for code auto-completion, encompassing a broad spectrum of algorithmic and data structural options.
Potential for application on LLM: The TACO dataset holds considerable promise for applications grounded in Large Language Models (LLMs). This dataset serves multiple functions: it can expedite the software development process through the automated generation of task-specific code samples, and it enhances the capabilities of LLMs in synthesizing code annotations and documentation to bolster both code comprehensibility and maintainability. Additionally, LLMs that leverage the TACO dataset are capable of addressing queries pertinent to programming and algorithmic topics, providing developers and learners with comprehensive and accurate responses. Collectively, these applications contribute to enhancements in programming efficiency, code quality, and the acquisition of programming knowledge, thereby stimulating ongoing advancements in the fields of programming and algorithms.